OpenStudy (anonymous):

Divide: 25xy4 + 60xy3 ------------- 5xy

OpenStudy (anonymous):

Distribute out 5xy from the top....

OpenStudy (anonymous):

what is distribute?

OpenStudy (anonymous):

a(X + Y) when you are asked in my example to distribute in a you have to use the foil method a(X + Y)= aX +aY

OpenStudy (anonymous):

If I ask you to distribute out a from aX +aY You do the opposite aX +aY= a( X + Y )